Solve : dy/dx=y 3 /(e 2x -y 2 )

Solve :
dy/dx=y3/(e2x-y2)

Grade:12

1 Answers

Yash Jain
9 years ago
Consider dx/dy
dx/dy = (e2x–y2)/y3
Put ex=t
gives, exdx/dy = dt/dy
So, dt/dy = t3/y3 – t/y
Now, let t=vy
so, dt/dy = v+ydv/dy  
gives, v+ydv/dy=v3-v
therefore, ∫dv/(v3-2v) = ∫dv/v(v+rt2)(v-rt2) ∫dy/y
Now you can solve it can’t you..
just by partial fractions.